OpenStudy (bookworm14):

Help Please rewrite the following formula :

OpenStudy (stefrheart):

okay..

OpenStudy (stefrheart):

where is it?

OpenStudy (anonymous):

\[\cal\large\color{blue}{Help~Please!!!}\color{red}{Rewrite~the~following~formula:}\]

OpenStudy (stefrheart):

how are you guys doing that?

OpenStudy (anonymous):

through the magic of \(\LaTeX\)

OpenStudy (bookworm14):

Woops I messed up me LaTeX lol ... but here is the prob for my math ----->

OpenStudy (anonymous):

divide both sides by \(4r^4\) and get \[t=\frac{S}{4r^4}\] in one step

OpenStudy (bookworm14):

@stefrheart its called \(\cal\color{purple}{LaTeX}\) its amazinggggg nad @satellite73 can you help me with possibly 1+ more questions?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

any one who knows latex and do these problems happy to help

OpenStudy (stefrheart):

ohh!

OpenStudy (bookworm14):

I'm lost on this one \(\huge\ddot\frown\)

OpenStudy (anonymous):

\[\huge h-m=h+fx\]

OpenStudy (anonymous):

subtract \(h\) from both sides first \[\huge -m=fx\]

OpenStudy (bookworm14):

we are supposed to solve for x did you see the answer choices?

OpenStudy (anonymous):

then divide by \(f\) and get\[\huge -\frac{m}{f}=x\]

OpenStudy (bookworm14):

are you for sure on these (please do not be offended)

OpenStudy (anonymous):

pretty darned sure check my work, see what you think

OpenStudy (bookworm14):

Okay fwewww ....okay here is another one (thank yo ufor helping) \(\Huge\color{teal}{\bigstar}\)

OpenStudy (bookworm14):

pshhh never mind I reread the problem I can do these last few on my own thank you for your help on the pther ones !! Wish me luck @satellite73 !!!!!!!!

OpenStudy (anonymous):

\[-4y+6=-3(y-7)\] distribute first

OpenStudy (anonymous):

k yw

OpenStudy (anonymous):

good luck !